# Galactica: A Scientific Large Language Model in Geoscience

### v1

@@ 3 Try
We have attempted to implement Megatron-LM [54] 292929https://github.com/NVIDIA/Megatron-LM, Megatron-Deepspeed 303030https://github.com/microsoft/Megatron-DeepSpeed and Huggingface-Deepspeed 313131https://huggingface.co/docs/transformers/main_classes/deepspeed approaches on the ROCm cluster with 2048 Hygon DCUs. We present a comparative analysis of the three techniques based on their distinctive features as follows:
@@
The magnitude of the data engineering and model training tasks would not have been possible without the collaborative efforts of multiple teams, specifically the data team, K2 team, architecture team, and model team from the Acemap323232https://www.acemap.info/ and LUMIA333333https://github.com/LUMIA-Group group in Shanghai Jiao Tong University and the team from the Institute of Geographical Science and Natural Resources Research, Chinese Academy of Sciences. The detailed contributions are as follows.
@@
GeoBench: Qiyuan Chen, Tianyu Huang, Cheng Deng.
@@
Inference Acceleration: Cheng Deng, Le Zhou.
Demo and API: Cheng Deng, Beiya Dai, Tianhang Zhang.
@@
Cheng Deng, Zhouhan Lin, and Yi Xu wrote the main paper, and Yutong Xu, Zhongmou He, Yuanyuan Shi, Yuncong Song, Tianhang Zhang, and Le Zhou wrote the Appendix.
@@
Chengdu University of Technology: Lei Zhang, Han Wang, Yangfan Liu, Tianyu Huang.
Nanjing Institute of Geology and Palaeontology, CAS: Yiwei Xu.

### v2

@@ Appendix I Appendix: Tool Learning Use cases
We have attempted to implement Megatron-LM [54] 323232https://github.com/NVIDIA/Megatron-LM, Megatron-Deepspeed 333333https://github.com/microsoft/Megatron-DeepSpeed and Huggingface-Deepspeed 343434https://huggingface.co/docs/transformers/main_classes/deepspeed approaches on the ROCm cluster with 2048 Hygon DCUs. We present a comparative analysis of the three techniques based on their distinctive features as follows:
@@
The magnitude of the data engineering and model training tasks would not have been possible without the collaborative efforts of multiple teams, specifically the data team, K2 team, architecture team, and model team from the Acemap353535https://www.acemap.info/ and LUMIA363636https://github.com/LUMIA-Group group in Shanghai Jiao Tong University and the team from the Institute of Geographical Science and Natural Resources Research, Chinese Academy of Sciences. The detailed contributions are as follows.
@@
GeoBench: Yuxun Miao, Qiyuan Chen, Cheng Deng.
@@
Inference Acceleration: Bo Xue, Cheng Deng, Le Zhou.
Demo and API: Bo Xue, Cheng Deng, Beiya Dai, Tianhang Zhang.
@@
Cheng Deng, Zhouhan Lin, and Yi Xu wrote the main paper, and Yutong Xu, Zhongmou He, Yuanyuan Shi, Yuncong Song, Tianhang Zhang, Bo Xue, and Le Zhou wrote the Appendix.
@@
Chengdu University of Technology: Lei Zhang, Han Wang, Yangfan Liu.